OpenStudy (anonymous):

65 × 35 + 50 × m = 4275 65 × 35 + 50 × m = 4275 @Mathematics

OpenStudy (anonymous):

how to solve m

OpenStudy (anonymous):

50m = 4275 - 2275 = 2000 m = 2000/50 = 40